OpenStudy (anonymous):

Will medal! Using a directrix of y = 2 and a focus of (3, -4), what quadratic function is created?

OpenStudy (anonymous):

@InstagramModel @phi @Directrix @AloneS @kiwibyrd anyone???

OpenStudy (anonymous):

These are the options: f(x) = 1/12 (x-3)^2 -1 f(x) = -1/6 (x+3)^2 +1 f(x) = 1/6 (x-3)^2 + 1 f(x) = -1/12 (x-3)^2 -1

OpenStudy (kiwibyrd):

uhhhh i dont know how to do these but if you go to mathway.com and type in all of the equations, you can check one by one to see which one matches the directrix and focus

OpenStudy (anonymous):

okay thanks for responding!

OpenStudy (therawbugeyes):

hey i can see if i can help ok

OpenStudy (anonymous):

thank you @therawbugeyes please

OpenStudy (therawbugeyes):

so do you know how to set this up?

OpenStudy (anonymous):

no, but i found the vertex?

OpenStudy (therawbugeyes):

ok and what did you get?

OpenStudy (anonymous):

(0, -1) if i did it correct

OpenStudy (therawbugeyes):

thats what i got

OpenStudy (anonymous):

where do i go from there?

OpenStudy (therawbugeyes):

here try this link it should help you ok http://jwilson.coe.uga.edu/emt668/EMAT6680.2002.Fall/Ledford/Ledford6/parabola.html

OpenStudy (therawbugeyes):

hey since your here can you see if you can help me?

OpenStudy (anonymous):

i can try

OpenStudy (therawbugeyes):

ok cool thxs:)

OpenStudy (therawbugeyes):

The equation below represents Function A and the graph represents Function B: Function A f(x) = 6x – 1 Function B graph of line going through ordered pairs 1, 4 and negative 1, negative 2 and negative 2, negative 5 Which equation best compares the slopes of the two functions? Slope of Function B = 2 x Slope of Function A. Slope of Function A = Slope of Function B Slope of Function A = 2 x Slope of Function B Slope of Function B = – Slope of Function A

OpenStudy (therawbugeyes):

did you get it?

OpenStudy (anonymous):

I tried but I've never seen anything like that @therawbugeyes sorry :/ not multiplying slopes